OpenStudy (anonymous):

use the Pythagorean theorem sqrt3^2 + 2sqrt3^2 = c^2 please help!!!

OpenStudy (hunus):

\[\sqrt{x^2}=x\] so \[\sqrt{3^2}=3\]

OpenStudy (anonymous):

the sqaure is not under the square root

OpenStudy (hunus):

\[(\sqrt{3})^2=3\] It's the same either way

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so it would be 3 + 2(3) = c^2 ?

OpenStudy (hunus):

Yes :)

OpenStudy (anonymous):

so then 9 would equal c^2?

OpenStudy (hunus):

Yup

OpenStudy (anonymous):

thank you!

OpenStudy (hunus):

No problem :)